OpenStudy (anonymous):

Media Services charges 20$ for a phone and $30 a month for its economy plan.The equation C=30+20 describs the totl cost,c, of operating a Media Services Phone for t months.The Total cost of ten months service is? If a customer had only 90 dollars how many months of service would this buy?

OpenStudy (anonymous):

I assume you meant C = 30t + 20

OpenStudy (anonymous):

yes, sorry.

OpenStudy (anonymous):

So for 10 months - substitute 10 in for t C = 30(10) + 20 C = 300 + 20 C = 320

OpenStudy (anonymous):

The equation should include T, as in: C = 30t + 20 Total cost for 10 months: \[\Large C=10*30+20\]\[\Large=320\] To find how long $90 would last, solve for t: \[\Large 90 = 30t + 20\]

OpenStudy (anonymous):

If you only have $90 you would put that into the total Cost 90 = 30t + 20 and solve for the number of months

OpenStudy (anonymous):

so the customer would get about 2.5 months service right?

OpenStudy (anonymous):

Around that: \[\Large \begin{array}{l} 90 = 30t + 20\\ 30t = 90 - 20\\ t = \frac{{70}}{{30}}\\ = \frac{7}{3} \end{array}\] Or about 2.333...

OpenStudy (anonymous):

yeah okay 2 months, 33 hours and 3 minutes of talk time. lol

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No worries :) Although it would actually be 1/3 of a MONTH... or about 10 days.

OpenStudy (anonymous):

I would say 2 months... I don't think they are going to give you a partial month.

OpenStudy (anonymous):

Yes blexting is right, given that's it monthly and you cannot exceed the $90 limit, you would have to stop at 2 months.

OpenStudy (anonymous):

2 months it is.